OpenStudy (anonymous):

Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2

OpenStudy (aaronq):

f(x) = −x + 4 -> f(g(x))= −(−2x−3) + 4 = 2x + 3 + 4 = 2x+7 f(g(2))= 2(2) +7

OpenStudy (anonymous):

g(2) = -2(2)-3 g(2) = -4-3 g(2) = -7 f(g(2)) =f(-7) f(-7) = -(-7)+4 f(-7) = 7+4 f(-7) = 11